# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 4.3. Verma modules
In this section, we investigate the objects in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ corresponding to the Verma modules under the equivalence ${\mathfrak{F}}^{b}$, and then show that they admit natural $B$-equivariant structures.
@@
Set $e=B/B\in{\mathcal{B}}$ and let $U_{e}=N^{-}B/B$ be the big open cell. By [44, Proof of Prop 3.12], there is an algebra isomorphism
@@
The counit $\epsilon:{\mathbb{C}}_{\zeta}[N^{-}]\rightarrow{\mathbb{C}}$ by evaluating at $1\in U^{\leq}_{\zeta}$ defines an ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ supported on $e$. As ${\mathcal{O}}_{{\mathcal{B}}}$-modules, ${\mathcal{O}}_{\epsilon}={\mathcal{O}}_{e}$.
@@
where ${\mathbb{C}}[T]\rightarrow Z^{\leq}_{\mathrm{Fr}}$ is given by $k_{\mu}\mapsto K^{2l}_{\mu}$, and ${\mathbb{C}}[T]\rightarrow{\mathbb{C}}[T]$ is given by $k_{\mu}\mapsto k^{l}_{\mu}$ (here we use $\{k_{\mu}\}_{\mu\in\Lambda}$ as the canonical basis of ${\mathbb{C}}[T]$). By [44, Lem 6.3], the map (4.10) induces an isomorphism of ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}$-modules
@@
We study the action of ${\mathfrak{U}}^{0}_{\zeta}$ on the LHS of (4.11). Fix basis $\{x_{r}\}_{r}$ in $U^{+}_{\zeta}$ and $\{y_{r}\}_{r}$ in ${\mathfrak{U}}^{-}_{\zeta}$ which are dual to each other under the canonical pairing between $U^{+}_{\zeta}$ and ${\mathfrak{U}}^{-}_{\zeta}$. By [44, Prop 4.13]  [^2] Because of our different conventions on the flag variety and the comultiplications of the quantum group, the formulas in loc. cit. are adapted after exchanging $E_{i}\leftrightarrow F_{i}$ and $K_{\lambda}\leftrightarrow K^{-1}_{\lambda}$. , on the LHS of (4.10), there is an equality for any dominant $\lambda\in\Lambda$,
@@
$\equiv\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}^{(1)}.1_{\lambda}\rangle\cdot y_{r}^{(2)}\cdot k_{\lambda}$
$=\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}.1_{\lambda}\rangle\cdot k_{\lambda}$
@@
where the third equality is because that $\Delta(y_{r})=y_{r}\otimes 1+\sum_{\text{deg}(y_{r}^{(2)})<0}y_{r}^{(1)}\otimes y_{r}^{(2)}$, and the last equality is because that the quantum Casimir element $\sum_{r}\mathrm{S}(x_{r})y_{r}$ acts on $1_{\lambda}\in{\mathrm{W}}(\lambda)$ by scalar $\zeta^{4(\rho,\lambda)}$. Finally, we identify
@@
Then we have ${\mathbb{M}}=\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}$ as ${\mathfrak{U}}^{\leq}_{\zeta}$-modules. It is in fact an isomorphism of ${\mathfrak{U}}_{\zeta}$-modules, since these two modules are $\Lambda$-graded. ∎
Recall the tilting bundle ${\mathcal{E}}$ in $\tilde{{\mathfrak{g}}}$. Let ${\mathcal{E}}^{\vee,b}$ be the restriction of the dual bundle ${\mathcal{E}}^{\vee}$ on $\{e\}\times{\mathfrak{b}}$ in $\tilde{{\mathfrak{g}}}$, and set ${\mathcal{E}}^{\vee,b}_{S}={\mathcal{E}}^{\vee,b}\otimes_{{\mathbb{C}}[{\mathfrak{t}}]}S$, which is an ${\mathrm{A}}^{b}_{S}$-module.
@@
As a sheave on ${\mathcal{V}}$, $\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon}$ is supported on the fiber of $e\in{\mathcal{B}}$ in ${\mathcal{V}}$, which is identified with $B\times_{T}T$ in the proof of the lemma above. We denote by $e_{n}$ the $n$-th infinitesimal neighborhood of 00 in ${\mathfrak{b}}$. We denote by $e^{\lambda}_{n}$ the $n$-th infinitesimal neighborhood of $(1,\zeta^{2\lambda})$ in $B\times_{T}T$, and view it as a closed subscheme in ${\mathcal{V}}$. Consider the subsheaf of algebras ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$ in $\widetilde{{\mathcal{D}}}_{\zeta}$. The ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ and the ${\mathcal{O}}_{\mathcal{V}}$-module ${\mathcal{O}}_{e^{\lambda}_{n}}$ define a sheaf of module ${\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$ of ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$. We abbreviate $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{\lambda}_{n})}=\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$. By Lemma 4.8 and formula (4.12), we compute that
@@
$R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{0},-):D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(1,1)}){\ \xrightarrow{\sim}\ }D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}).$
@@
$={\mathfrak{M}}_{0}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}.$
@@
${\mathfrak{M}}_{0}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}={\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$
@@
${\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$ $=R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})$
@@
By the equivalence (4.6), the RHS of (4.14) is equal to ${\mathcal{O}}_{e_{n}}$ (note that we identify the formal neighborhood of 00 in ${\mathfrak{b}}_{S}$ with the formal neighborhood of $(1,1)$ in $B\times_{T}T$).
Finally, under the equivalence (3.1), the sheaf ${\mathcal{O}}_{e_{n}}$ corresponds to the ${\mathrm{A}}$-module ${\mathcal{E}}^{\vee}\otimes_{{\mathcal{O}}_{\tilde{{\mathfrak{g}}}}}{\mathcal{O}}_{e_{n}}={\mathcal{E}}^{\vee,b}_{S}/{\mathfrak{m}}^{n}_{0}{\mathcal{E}}^{\vee,b}_{S}$, where ${\mathfrak{m}}_{0}$ is the maximal ideal of $0\in{\mathfrak{b}}$ in ${\mathbb{C}}[{\mathfrak{b}}]$. Hence we have
@@
We define the Verma objects in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ to be
@@
An object in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}$ admits Verma flags if it admits a (decreasing, separated, ${\mathbb{Z}}_{\leq 0}$-) filtration with composition factors given by Verma objects.
@@
Any Verma object in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ admits a natural $B$-equivariant structure.
@@
By Proposition 4.9, ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ is naturally a $B$-equivariant ${\mathrm{A}}^{b}_{S}$-module. We show that in general
(4.15) ${\mathcal{M}}_{t(\nu)w_{0}w^{-1}}={\mathbb{T}}^{c}_{w}({\mathcal{M}}_{w_{0}})\otimes{\mathbb{C}}_{\nu},\quad\forall w\in W,\ \forall\nu\in\Lambda,$
where ${\mathbb{T}}^{c}_{w}={\mathbb{T}}^{c}_{s_{1}}\cdots{\mathbb{T}}^{c}_{s_{n}}$ for a reduced expression $w=s_{1}\cdots s_{n}$. Then since ${\mathbb{T}}^{c}_{w}$ is $B$-equivariant, any Verma object is endowed with a $B$-equivariant structure.
Indeed, for any $w\in W$ and $s\in{\mathrm{I}}$ such that $ws<w$ in the Bruhat order, there is a short exact sequence (see e.g. [23, Thm 7.14(a)])
[…]
@@
Apply (4.16) several times on ${\mathcal{M}}_{w_{0}}$ and use the $\Lambda$-translations, we obtain (4.15). ∎
4.4. Fully-faithfulness
@@
as in §4.2.2. Let ${\mathbb{K}}={\mathrm{Frac}}(S)$ be the fraction field of $S$.
@@
$\mathrm{Hom}({\mathbb{V}}^{b}_{r,R}M_{1},{\mathbb{V}}^{b}_{r,R}M_{2})\hookrightarrow\mathrm{Hom}(T^{-\rho}_{0}M_{1},T^{-\rho}_{0}M_{2}),\quad\forall M_{i}.$
Hence the faithfulness of ${\mathbb{V}}^{b}_{r,R}$ (on objects admitting Verma flags) follows from the faithfulness of $T^{-\rho}_{0}$, see [36, Lem 4.9(2)].
@@
${\mathbb{V}}^{b}_{r,{\mathbb{K}}}=\bigoplus_{w\in W}T^{-\rho}_{0}\circ\iota_{w},$
where $\iota_{w}$ is the inclusion from the block labelled by $w$ in (4.19). The functor $T^{-\rho}_{0}\circ\iota_{w}$ is an equivalence with quasi-inverse given by $\mathrm{pr}_{w}\circ T^{0}_{-\rho}$, where $\mathrm{pr}_{w}$ is the corresponding block projection. Hence ${\mathbb{V}}^{b}_{r,{\mathbb{K}}}$ is an equivalence of categories. ∎
@@
By Corollary 4.7, we only have to prove the assertion for ${\mathbb{V}}^{b}_{r}$. It follows standardly from Lemma 4.12 by the general theory of highest weight categories, see e.g. [35, Prop 2.18]. For readers’ convenience, we present the proof as follows.
We abbreviate ${\mathbb{V}}={\mathbb{V}}^{b}_{r,R}$ and $M_{R}=M\otimes_{S}R$ for any $S$-module $M$ and any $S$-algebra $R$. Let $M_{1},M_{2}$ be objects in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ admitting Verma flags. In particular they are free as $S$-modules. Hence the base change yields an embedding $\mathrm{Hom}(M_{1},M_{2})\subseteq\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$. By [32, Thm 38], we have $M_{i}=\bigcap\limits_{\mathrm{ht}(\mathfrak{p})=1}M_{i,S_{\mathfrak{p}}}$ as subspaces in $M_{i,{\mathbb{K}}}$ (the intersection is taken for height 1 prime ideals $\mathfrak{p}$ of $S$). Hence we can identify
@@
00{0}${\mathrm{Hom}(M,M^{\prime})_{R^{\prime}}}$${\mathrm{Hom}(P_{2},M^{\prime})_{R^{\prime}}}$${\mathrm{Hom}(P_{1},M^{\prime})_{R^{\prime}}}$00{0}${\mathrm{Hom}(M_{R^{\prime}},M^{\prime}_{R^{\prime}})}$${\mathrm{Hom}(P_{2,R^{\prime}},M^{\prime}_{R^{\prime}})}$${\mathrm{Hom}(P_{1,R^{\prime}},M^{\prime}_{R^{\prime}}),}$
where the subscript $R^{\prime}$ means $-\otimes_{R}R^{\prime}$. By [20, Prop 2.4], the right two vertical maps are isomorphisms, so is the left one. ∎
@@
[…]
where the vertical isomorphisms are by Lemma 4.14 applied on the tuple $(M,M^{\prime},R,R^{\prime})$ $=(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}},S_{\mathfrak{p}},{\mathbb{K}})$ or $({\mathbb{V}}M_{1,S_{\mathfrak{p}}}$, ${\mathbb{V}}M_{2,S_{\mathfrak{p}}}$, $S_{\mathfrak{p}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$, ${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}])$, and the lower horizontal isomorphism is by Lemma 4.12(2). Hence $\varphi\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$ is an isomorphism. We conclude that $\varphi$ is an isomorphism. ∎
@@
For the non-deformed category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}$, the functor ${\mathbb{V}}^{b}_{r,{\mathbb{C}}}$ fails to be full-faithful on objects admitting Verma flags.

### v4

@@ 4.5. Verma modules
In this section, we investigate the objects in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ corresponding to the Verma modules under the equivalence ${\mathfrak{F}}^{b}$, and then show that they admit natural $B$-equivariant structures.
@@
Set $e=B/B\in{\mathcal{B}}$ and let $U_{e}=N^{-}B/B$ be the big open cell. By [42, Proof of Prop 3.12], there is an algebra isomorphism
@@
[…]
@@
where ${\mathbb{C}}[T]\rightarrow Z^{\leq}_{\mathrm{Fr}}$ is given by $k_{\mu}\mapsto K^{2l}_{\mu}$, and ${\mathbb{C}}[T]\rightarrow{\mathbb{C}}[T]$ is given by $k_{\mu}\mapsto k^{l}_{\mu}$ (here we use $\{k_{\mu}\}_{\mu\in\Lambda}$ as the canonical basis of ${\mathbb{C}}[T]$). By [42, Lem 6.3], the map (4.16) induces an isomorphism of ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}$-modules
@@
We study the action of ${\mathfrak{U}}^{0}_{\zeta}$ on the LHS of (4.17). Fix basis $\{x_{r}\}_{r}$ in $U^{+}_{\zeta}$ and $\{y_{r}\}_{r}$ in ${\mathfrak{U}}^{-}_{\zeta}$ which are dual to each other under the canonical pairing between $U^{+}_{\zeta}$ and ${\mathfrak{U}}^{-}_{\zeta}$. By [42, Prop 4.13]  [^1] Because of our different conventions on the flag variety and the comultiplications of the quantum group, the formulas in loc. cit. are adapted after exchanging $E_{i}\leftrightarrow F_{i}$ and $K_{\lambda}\leftrightarrow K^{-1}_{\lambda}$. , on the LHS of (4.16), there is an equality for any dominant $\lambda\in\Lambda$,
@@
$\equiv\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}^{(1)}1_{\lambda}\rangle\cdot y_{r}^{(2)}\cdot k_{\lambda}$
$=\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}1_{\lambda}\rangle\cdot k_{\lambda}$
@@
where the third equality is because $\Delta(y_{r})=y_{r}\otimes 1+\sum_{\text{deg}(y_{r}^{(2)})<0}y_{r}^{(1)}\otimes y_{r}^{(2)}$, and the last equality is because the quantum Casimir element $\sum_{r}\mathrm{S}(x_{r})y_{r}$ acts on $1_{\lambda}\in{\mathrm{W}}(\lambda)$ by scalar $\zeta^{4(\rho,\lambda)}$. Finally, we identify
@@
Then we have ${\mathbb{M}}=\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}$ as ${\mathfrak{U}}^{\leq}_{\zeta}$-modules. It is in fact an isomorphism of modules of ${\mathfrak{U}}_{\zeta}$, since the two modules are $\Lambda$-graded. ∎
Recall the tilting bundle $\mathcal{E}$} in $\widetilde{{\mathfrak{g}}}$. Let ${\mathcal{E}}^{\vee,b}$ be the restriction of the dual bundle $\mathcal{E}$$^{\vee}$ on $\{e\}\times{\mathfrak{b}}$ in $\widetilde{{\mathfrak{g}}}$, and set ${\mathcal{E}}^{\vee,b}_{S}={\mathcal{E}}^{\vee,b}\otimes_{{\mathbb{C}}[{\mathfrak{t}}]}S$, which is an ${\mathsf{A}}^{b}_{S}$-module.
@@
As a sheave on $\mathcal{V}$}, $\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon}$ is supported on the fiber of $e\in{\mathcal{B}}$ in $\mathcal{V}$}, which is identified with $B\times_{T}T$ in the proof of the lemma above. We denote by $e_{n}$ the $n$-th infinitesimal neighborhood of 0 in $\mathfrak{b}$}. We denote by $e^{\lambda}_{n}$ the $n$-th infinitesimal neighborhood of $(1,\zeta^{2\lambda})$ in $B\times_{T}T$, and view it as a closed subscheme in $\mathcal{V}$}. Consider the subsheaf of algebras ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$ in $\widetilde{{\mathcal{D}}}_{\zeta}$. The ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ and the ${\mathcal{O}}_{\mathcal{V}}$-module ${\mathcal{O}}_{e^{\lambda}_{n}}$ define a sheaf of module ${\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$ of ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$. We abbreviate $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{\lambda}_{n})}=\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$. By Lemma 4.8 and formula (4.18), we compute that
@@
[…]
@@
$={\mathfrak{M}}_{0}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}.$
@@
${\mathfrak{M}}_{0}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}={\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$
@@
${\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$ $=R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})$
@@
By the equivalence (4.6), the RHS of (4.20) is equal to ${\mathcal{O}}_{e_{n}}$ (note that we identify the formal neighborhood of 0 in ${\mathfrak{b}}_{S}$ with the formal neighborhood of $(1,1)$ in $B\times_{T}T$).
Finally, under the equivalence (3.1), the sheaf ${\mathcal{O}}_{e_{n}}$ corresponds to the $\mathsf{A}$}-module ${\mathcal{E}}^{\vee}\otimes_{{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}}}{\mathcal{O}}_{e_{n}}={\mathcal{E}}^{\vee,b}_{S}/{\mathfrak{m}}^{n}_{0}{\mathcal{E}}^{\vee,b}_{S}$, where ${\mathfrak{m}}_{0}$ is the maximal ideal of $0\in{\mathfrak{b}}$ in ${\mathbb{C}}[{\mathfrak{b}}]$. Hence we have
@@
We define the Verma objects in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ to be
@@
An object in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}$ admits Verma flags if it admits a filtration of finite length with composition factors given by Verma objects.
For any $w\in W$, we choose a reduced expression $w=s_{1}\cdots s_{n}$ and set ${\mathbb{T}}^{c}_{w}={\mathbb{T}}^{c}_{s_{1}}\cdots{\mathbb{T}}^{c}_{s_{n}}$.
@@
For any $w\in W$ and $\nu\in\Lambda$, we have
(4.21) ${\mathcal{M}}_{t(\nu)w_{0}w^{-1}}={\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee,b}_{S})\otimes{\mathbb{C}}_{\nu},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
As a consequence, any Verma object in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ admits a natural $B$-equivariant structure.
@@
For any $w\in W$ and $s\in{\mathrm{I}}$ such that $ws<w$ in the Bruhat order, there is a short exact sequence (see e.g. [23, Thm 7.14(a)])
$0\rightarrow M(ws\bullet 0)_{S}\rightarrow\Theta^{r}_{s}(M(w\bullet 0))_{S}\xrightarrow{\text{counit}}M(w\bullet 0)_{S}\rightarrow 0,$
@@
Applying (4.22) several times on ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ (Proposition 4.9) and using the $\Lambda$-translations, we obtain (4.21). Finally, since ${\mathcal{E}}^{\vee,b}_{S}$ is naturally endowed with a $B$-equivariant structure and ${\mathbb{T}}^{c}_{w}$ is $B$-equivariant, any Verma object is $B$-equivariant. ∎
4.6. Fully-faithfulness for ${\mathbb{V}}^{b}$
@@
as in §4.4 (note that for $S=R$ we abbreviate ${\mathbb{V}}^{b}_{r}={\mathbb{V}}^{b}_{r,S}$). Let ${\mathbb{K}}={\mathrm{Frac}}(S)$ be the fraction field of $S$.
@@
$\mathrm{Hom}({\mathbb{V}}^{b}_{r,R}M_{1},{\mathbb{V}}^{b}_{r,R}M_{2})\hookrightarrow\mathrm{Hom}({\mathsf{T}}^{-\rho}_{0}M_{1},{\mathsf{T}}^{-\rho}_{0}M_{2}),\quad\forall M_{i}.$
Hence the faithfulness of ${\mathbb{V}}^{b}_{r,R}$ (on objects admitting Verma flags) follows from the faithfulness of ${\mathsf{T}}^{-\rho}_{0}$, see [34, Lem 4.9(2)].
@@
${\mathbb{V}}^{b}_{r,{\mathbb{K}}}=\bigoplus_{w\in W}{\mathsf{T}}^{-\rho}_{0}\circ\iota_{w},$
where $\iota_{w}$ is the inclusion from the block labelled by $w$ in (4.25). The functor ${\mathsf{T}}^{-\rho}_{0}\circ\iota_{w}$ is an equivalence with quasi-inverse given by $\mathrm{pr}_{w}\circ{\mathsf{T}}^{0}_{-\rho}$, where $\mathrm{pr}_{w}$ is the corresponding block projection. Hence ${\mathbb{V}}^{b}_{r,{\mathbb{K}}}$ is an equivalence of categories. ∎
@@
By (4.15) we only have to prove the assertion for ${\mathbb{V}}^{b}_{r}$. It follows standardly from Lemma 4.12 by the general theory of highest weight categories, see e.g. [33, Prop 2.18]. For readers’ convenience, we present the proof as follows.
We abbreviate ${\mathbb{V}}={\mathbb{V}}^{b}_{r,R}$ and $M_{R}=M\otimes_{S}R$ for any $S$-module $M$ and any $S$-algebra $R$. Let $M_{1},M_{2}$ be objects in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ admitting Verma flags. In particular they are free as $S$-modules. Hence the base change yields an embedding $\mathrm{Hom}(M_{1},M_{2})\subseteq\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$. By [30, Thm 38], we have $M_{i}=\bigcap\limits_{\mathrm{ht}(\mathfrak{p})=1}M_{i,S_{\mathfrak{p}}}$ as subspaces in $M_{i,{\mathbb{K}}}$ (the intersection is taken for height 1 prime ideals $\mathfrak{p}$ of $S$). Hence we can identify
@@
[…]
where the subscript $R^{\prime}$ means $\otimes_{R}R^{\prime}$. By [20, Prop 2.4], the right two vertical maps are isomorphisms, so is the left one. ∎
@@
$\mathrm{Hom}(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}})\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$${\mathrm{Hom}({\mathbb{V}}M_{1,S_{\mathfrak{p}}},{\mathbb{V}}M_{2,S_{\mathfrak{p}}})}\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$$\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$$\mathrm{Hom}({\mathbb{V}}M_{1,{\mathbb{K}}},{\mathbb{V}}M_{2,{\mathbb{K}}}).$$\scriptstyle{\varphi\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$
where the vertical isomorphisms are by Lemma 4.14 applied on the tuple $(M,M^{\prime},R,R^{\prime})$ $(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}},S_{\mathfrak{p}},{\mathbb{K}})$ or $({\mathbb{V}}M_{1,S_{\mathfrak{p}}}$, ${\mathbb{V}}M_{2,S_{\mathfrak{p}}}$, $S_{\mathfrak{p}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$, ${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}])$, and the lower horizontal isomorphism is by Lemma 4.12(2). Hence $\varphi\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$ is an isomorphism. We conclude that $\varphi$ is an isomorphism. ∎
@@
For the non-deformed category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{\mathbb{C}}$, the functor ${\mathbb{V}}^{b}_{r,{\mathbb{C}}}$ fails to be fully-faithful on objects admitting Verma flags.